OpenStudy (anonymous):

Find dy/dx, using implicit differentiation the equation is 2x + 9y= xy

OpenStudy (matt101):

\[2x+9y=xy\]\[2+9y'=y+xy'\]\[9y'-xy'=y-2\]\[y'(9-x)=y-2\]\[y'=(y-2)/(9-x)\]

OpenStudy (anonymous):

what happens to the x at the veyr beginning?

OpenStudy (anonymous):

very*

OpenStudy (anonymous):

f(x)=y that's why when y(which is originally f(x) or x) is multiplied to y1 the y(which is originally f(x) or x) got cancelled.